# Measuring Round-Trip Response Latencies Under Asymmetric Routing

### v1

@@ Abstract
Latency is a key indicator of Internet service performance. Continuously tracking the latency of client requests enables service operators to quickly identify bottlenecks, perform adaptive resource allocation or routing, and mitigate attacks. Passively measuring the response latency at intermediate vantage points is attractive since it provides insight into the experience of real clients without requiring client instrumentation or incurring probing overheads.
@@
We argue that existing passive measurement techniques have not caught up with recent trends in service deployments, specifically, the increasing uptake of encrypted transports (such as QUIC) and the use of asymmetric routing by design. Existing methods are inapplicable, inaccurate, or inefficient.
@@
This paper presents Pirate, a passive approach to measure response latencies when only the client-to-server traffic is visible, even when transport headers are encrypted. Pirate estimates the time gap between causal pairs—two requests such that the response to the first triggered the second—as a proxy for the client-side response latency. Our experiments with a realistic web application show that Pirate can estimate the response latencies measured at the client application layer to within 1%. A Pirate-enhanced layer-4 load balancer (with DSR) cuts tail latencies by 37%.
1 Introduction
@@
Latency is a key indicator of the performance and quality of interactive Internet services. For developers of such services, it is well known that smaller client-visible latencies drive better user engagement [4, 3, 74, 14, 10]. Given its primacy, accurate measurements of latency can feed important decisions in designing and adapting networked systems. For example, operators of some large content delivery services use the latency between content servers and users to determine which servers to redirect users to [76, 103]. Autonomous Systems (ASes) on the wide-area Internet may use high or variable latencies experienced by transiting connections to identify pathologies such as persistent link congestion [49, 101, 47, 100] or interdomain route hijack [7, 18, 38], and take corrective actions to fix routing configurations or provision capacity [80]. Within a data center, server latency may be used to implement performance-optimized replica selection in load balancers [39, 57] or remote procedure call (RPC) clients [111, 102].
@@
Continuous measurement of client-visible latency is crucial since the latency may change over the lifetime of a connection [63, 36, 50, 40]. For example, time-varying bursts and packet losses in the network [59, 43], or server variability due to noisy neighbors, load, and resource scheduling [81, 56, 42, 40], can significantly change the latency perceived by the same client connection over time.
@@
Consequently, the community has developed several techniques and systems for continuous latency measurement. Broadly, active approaches send explicit probes that observe latency, for example, by running ICMP pings between servers [62]. Passive approaches observe latency at strategic locations that can be controlled [115, 103, 113], possibly at intermediate vantage points outside the client and the server. Passive approaches do not require instrumentation or self-reporting from clients which may be untrustworthy or not easily changed [103]. If passive measurement is possible, it can observe real and representative client connections [40], avoiding the typical downsides of active approaches that incur compute and network resources for probing [46, 75].
@@
This Paper. We are interested in passively and continuously measuring response latencies, which we define as the time between when an application-layer request is sent and the last byte of the response is delivered to the client application, i.e. the per-object time to last byte. For RPCs within data centers, response latency corresponds to the RPC completion time at the application layer [79, 68]. For web-based applications, the response latencies of specific objects (e.g. those first painted or largest on the user’s screen) are highly correlated with several quality of experience metrics [41, 29, 30, 21]. As we elaborate in §2.1, to be applicable to a broad range of scenarios, we seek techniques to passively measure response latency while meeting the following additional requirements:
  1. 1.
    handle encrypted transport and application headers;
  2. 2.
@@
  3. 3.
    generalize across transports and transport algorithms; and
  4. 4.
@@
At first glance, it may appear that the response latency is closely related to the transport-layer round-trip time (RTT), whose passive measurement is widely studied, e.g. [96, 60, 95, 107, 66, 45, 50, 67]. However, response latency is distinct from the transport-layer RTT, since the server may return a transport-layer acknowledgment well before the full application-layer response. As far as we are aware, no existing RTT measurement technique can meet the additional requirements above, and we know of no passive techniques to directly measure response latency (§2.2).
@@
Our first idea is to leverage the closed-loop nature of Internet protocols and applications. When application requests depend on the contents of prior responses—for example, a web object embeds other objects—the reception of a prior response generates the subsequent request. Flow-controlled applications (transports) mandate that new requests (packets) are only transmitted when prior responses (acknowledgments) arrive. Such closed-loop packet transmission behavior enables estimating the response latency by proxy: the vantage point can measure the time delay between a request and a subsequent request that was triggered by the reception of the response to the first request. We call the latter request a causally-triggered request, and the pair of requests a causal pair. Many latency-sensitive applications exhibit cross-request dependencies (e.g. web [86], RPC [98, 82, 56, 23]) and flow control (e.g. key-value store [88], web [9, 15]).
@@
It is not obvious how to identify causal pairs. At any given time, many concurrent requests may be in flight. Connection persistence and stream multiplexing is standardized and widely deployed in the HTTP protocols [8, 19, 1]. Two consecutive requests arriving at a vantage point from the same connection do not necessarily form a causal pair.
@@
Our second key idea is to leverage the time gaps between packet arrivals to identify causal pairs. Once a client has sent a few requests, its transmission of subsequent requests is typically blocked by dependencies or flow control. Hence, causally-triggered requests arrive after a time gap that is longer than the inter-arrival times of the packets just prior. We call this the prominent packet gap assumption. By choosing an appropriate threshold on packet inter-arrival times, a vantage point can classify any packet arriving with a time gap exceeding this threshold as a causally-triggered request. The time interval between two consecutive causally-triggered requests provides one estimate of the response latency.
@@
But what should the time threshold be set to? How can we make this setting robust across deployments (e.g. wide area vs data center), transports (e.g. TCP vs QUIC), and applications (e.g. web vs key-value store)?
@@
Our third key idea is the following: While a single prominent packet gap is subject to (unknown) network, application, and server conditions, observing packet inter-arrivals over a period of time provides a more robust picture. In this paper, we leverage the probability distribution of inter-arrival times to identify prominent packet gaps, by designing a lightweight construction to measure a coarse histogram of all packet inter-arrival times within a connection. We also devise a procedure to estimate the average response latency over the (configurable) time period when the distribution was measured.
@@
We call our algorithm, a synthesis of the three key ideas above, Pirate. We show how to use measurements from Pirate in a feedback control loop (§4), by adapting a layer-4 load balancer to leverage real-time response latencies. This design unilaterally changes the load balancer, leaving clients, servers, application software, and the network unmodified. The load balancer may use direct server return [90] to avoid processing the responses.
Results. In §5, we evaluate Pirate under a realistic web workload derived from the Alexa top-100 web sites, with a web server whose CPU availability varies follows a real CPU utilization trace [108]. Across all monitored responses, Pirate achieves a median relative error of 0.63% relative to the response latency measured at the client application. Our results affirm the hypothesis that the proxy request-to-request intervals estimated by Pirate are sound approximations of client-application-visible response latencies. We have localized the causes of errors in Pirate to two main reasons: client-side processing delays and staggered flights of response packets transmitted by servers. Our results further show that transport-layer RTT measurement cannot model the application-layer response latency accurately, even when such measurement can take advantage of bidirectional visibility into traffic (Pirate only sees client-to-server traffic).
We integrated Pirate into Katran [11], an open-source layer-4 load balancer based on Maglev hashing [53]. Latency-aware Katran cuts the 99th percentile latency in our experimental setup by 37% on average across loads, relative to (unmodified) Katran. Latency-aware Katran also shows smaller variability in tail latencies. The accuracy of Pirate is robust to packet loss and reordering in the network. We also show that Pirate accurately estimates response latencies over an encrypted transport (mvfst QUIC [31]). Pirate imposes an average delay of 346 ns in the critical packet-processing path when forwarding packets using XDP [65], a fast packet-processing platform. For context, the Katran load balancer takes on average 1114 ns to process each packet.
@@
A shorter version of this paper appeared previously at a workshop venue (citation elided for anonymity). In comparison, this paper incorporates new techniques to make estimation more robust, significantly expands evaluations of accuracy and overhead under realistic settings, and designs a feedback controller that operates on real-time measurements.
2 Motivation and Background
@@
Monitoring latency is a fundamental requirement for the design, maintenance, and optimization of interactive Internet services. This paper studies the problem of measuring response latency, which we define as the time between when an application client sends out the request and when it receives the last byte of the corresponding response. The discussions below apply either when the client is a user device contacting a web server or an RPC client in one tier of a distributed multi-tiered application.
2.1 Our Goals
@@
We seek a passive measurement approach to continuously measure response latency at a vantage point which can lie outside the client and the server, but along the path from the client to the server. To make it more broadly applicable, we also seek to meet the following more specific goals.
@@
G1. Handle encrypted transport headers. Encryption of application-layer payloads is ubiquitous on the web and in compute clusters [20]. Modern transport protocols and network-layer security go one level further, obscuring transport-layer headers. Deployments of HTTP/3 over QUIC [24, 16] and network-layer encrypted tunnels [34], both of which encrypt transport headers, are trending upward. We desire a method that does not require visibility into transport layer headers.
@@
G2. Handle routing asymmetry. Nodes conducting passive measurement may be at locations which do not have access to both directions of traffic flowing between the client and the server. The prevalence of routing asymmetry is well established in the wide-area Internet [91] and also in data centers [114]. Further, many network deployments use asymmetric routing by design. For example, layer-4 load balancers employ direct server return [90, 53]—bypassing the load balancer for response packets—to avoid redundant packet processing at the load balancer. BGP policies on the Internet (e.g. stub networks) may choose different ingress and egress border routers for a given connection [71, 104].
@@
G3. Generalize across transports and transport algorithms. The design of new TCP and QUIC algorithms (e.g. congestion control, loss recovery) is an active and rapidly evolving area. We seek techniques that avoid relying on specific behaviors of the transport protocol, generalize across transports and transport algorithms, and sidestep the need for new extensions to transport protocols.
@@
G4. Capable of being run “online,” in particular on software middleboxes. Measurement and monitoring devices are frequently deployed in the form of virtual network functions managed through a software-defined measurement infrastructure, e.g. [89, 35]. We aim for techniques that impose low compute and memory overheads in software deployments, e.g. middleboxes running high-speed packet processing.
2.2 Prior Work and Its Applicability
Measuring latency is a deeply studied problem in networking. In particular, significant prior work exists on passive measurement to observe round-trip times (RTTs) at the transport layer. Below, we survey the literature and the suitability of the ideas therein to attaining our measurement goals (§2.1).
@@ 2.2 Prior Work and Its Applicability
RTT estimation approaches. The basic idea of passive RTT estimation is relating data packets to their corresponding transport-layer acknowledgments (ACKs) and measuring the time difference between this pair of packets. Many prior works on RTT estimation determine this relationship by connecting the sequence numbers or timestamps on data packets to the corresponding information echoed on ACK packets [96, 60, 95, 50, 107, 66]. Implementing this basic approach requires visibility into transport headers in the clear (G1), and seeing packets in both directions (G2).
@@
We are aware of only a small number of techniques which apply when the passive observor sees packet traffic in only one direction. An early approach from Jiang and Dovrolis [67] estimates RTT through the time interval between the SYN and the ACK packet of the 3-way handshake, and also between small packet bursts during the early rounds of slow start. These techniques are customized to the specific dynamics of the TCP protocol (G3) and do not measure connections beyond the beginning. Another set of prior works estimate RTTs by computing frequency spectra over the time series of packet arrivals [45, 107]. Such computations require significant memory to hold packet timings (G4), and can be complex to tune [50]. Most recently, researchers have proposed the spin bit as a protocol extension (G3) to passively measure RTTs [48]. It has been incorporated as an optional mechanism in QUIC [12], but its deployment may be limited [77].
@@
Response latency is dissimilar to the RTT. A high RTT implies a high response latency, but a low RTT does not necessarily imply a low response latency. In principle, the response latency can significantly differ from the transport-layer RTT. The server can transmit a transport-layer ACK to a client prior to transmitting any response data. Further, responses, typically larger than requests, may span multiple packets, widening the gap between the transport-layer ACK and the last byte of the response. When the client and server use a protocol with pipelined requests (e.g. HTTP/1.1, HTTP/2, and QUIC), multiple application requests may be transmitted in a single packet from client to server. The transport-layer ACK only corresponds to the first response.
Figure 1: TCP Round-Trip Time (RTT) estimation from the network stack (tcp-probe), trace-based methods (tcpTrace), and syn-ack estimation [67] do not match the client-visible response latency.
@@
Experimentally, we find that RTT and response latency differ even when there is only one small application-layer request in flight. We ran a web benchmark where the clients maintain a single outstanding request at a time with the server (our full experimental setup is described in §5.1). For one of the client threads, Figure 1 shows the time series of response latencies (marked req-to-res) against three techniques to measure RTT: (i) using the kernel network stack through tcp-probe [27]; (ii) by estimation from a packet trace (tcptrace [96]; and (iii) syn-ack estimation [67]. Not only is the response latency different from the RTT, but it is also uncorrelated.
@@
Other related work. Given the emerging prevalence of encrypted transports, some recent work aims to passively infer quality of experience metrics specifically for web applications [64, 109]. These metrics compose the response latencies of multiple objects, e.g. to estimate overall web experience [32], through machine learning and prediction. In contrast, we are concerned with continuous measurement of response latencies of individual objects throughout the lifetime of a connection. There is also significant prior work on passively measuring the latency of specific segments of a network, e.g. [75, 78, 59, 43] and algorithms to estimate latencies over end-to-end network paths using previously measured delays, e.g. [87, 83]. Plenty of active approaches exist to measure end-to-end network latency characteristics, e.g. [100, 112, 101]. This paper seeks continuous passive measurement techniques for end-to-end response latency.
@@
In summary, we believe that measuring response latencies passively and continuously under modern transports and practical deployment constraints (§2.1) is as yet an unsolved problem. In the next section, we discuss a measurement approach that addresses this problem.
3 Pirate

### v2

# Measuring Response Latencies Under Asymmetric Routing

@@ Abstract.
Latency is a key indicator of Internet service performance. Continuously tracking the latency of client requests enables service operators to quickly identify bottlenecks, perform adaptive resource allocation or routing, and mitigate attacks. Passively measuring response latency at intermediate vantage points is attractive because it provides insight into the experience of real clients without requiring client instrumentation or incurring probing overhead.
@@
We argue that existing passive measurement techniques have not caught up with recent trends in service deployments, specifically the increasing prevalence of obscured or encrypted transport headers and the use of asymmetric routing by design. Existing methods are inapplicable, inaccurate, or inefficient.
@@
This paper presents Pirate, a passive approach for measuring response latencies when only client-to-server traffic is visible, even when transport headers are invisible. Pirate estimates the time gap between causal pairs—two requests such that the response to the first triggered the second—as a proxy for the client-side response latency. Our experiments with a realistic web application show that Pirate can accurately estimate the response latencies measured at the client application layer. A Pirate-enhanced layer-4 load balancer (with DSR) cuts tail latencies by 37%.
copyright: noneconference: ; 2026
1. Introduction
@@
Latency is a key indicator of the performance and quality of interactive Internet services. For developers of such services, it is well known that smaller client-visible latencies drive better user engagement (114; 75; R. Kohavi, R. M. Henne, and D. Sommerfield (2007); 77; 80). Given its primacy, accurate latency measurements can feed important decisions in the design and adaptation of networked systems. For example, operators of some large content delivery services use the latency between content servers and users to determine which servers to redirect users to (Krishnan et al., 2009; Szymaniak et al., 2008). Autonomous Systems (ASes) on the wide-area Internet may use high or variable latencies experienced by transiting connections to identify pathologies such as persistent link congestion (Dhamdhere et al., 2018; Sundaresan et al., 2017; Choi et al., 2007; Sommers et al., 2007) or interdomain route hijack (107; 125; A. Arnbak and S. Goldberg (2015)), and take corrective actions to fix routing configurations or provision capacity (Lin et al., 2025). Within a data center, server latency may be used to implement performance-optimized replica selection in load balancers (Barbette et al., 2020; Gandhi and Narayana, 2025) or remote procedure call (RPC) clients (Wydrowski et al., 2024; Suresh et al., 2015).
@@
Continuous measurement of client-visible latency is crucial, as latency may change over the lifetime of a connection (Høiland-Jørgensen et al., 2016; Aikat et al., 2003; Ding and Rabinovich, 2015; Barford and Crovella, 1999). For example, time-varying bursts and packet losses in the network (Ghabashneh et al., 2022; Canel et al., 2024), or server variability due to noisy neighbors, load, and resource scheduling (Lo et al., 2015; Gan et al., 2019; Cai et al., 2021; Barford and Crovella, 1999), can significantly change the latency perceived by the same client connection over time.
@@
Consequently, the community has developed several techniques and systems for continuous latency measurement. Broadly, active approaches send explicit probes to observe latency, for example, by running ICMP pings between servers (Guo et al., 2015). Passive approaches observe latency at strategic locations that can be controlled (Zhu et al., 2015; Szymaniak et al., 2008; Yu et al., 2011), possibly at intermediate vantage points outside the client and the server. Passive approaches do not require instrumentation or self-reporting from clients, which may be untrustworthy or difficult to change (Szymaniak et al., 2008). If passive measurement is possible, it can observe real and representative client connections (Barford and Crovella, 1999), avoiding the typical downsides of active approaches that incur compute and network resources for probing (Choffnes and Bustamante, 2009; Kompella et al., 2009).
@@
This Paper. We are interested in passively and continuously measuring response latencies, which we define as the time between when an application-layer request is sent and the last byte of the response is delivered to the client application, i.e., the per-object time-to-last-byte. For RPCs within data centers, response latency corresponds to the RPC completion time at the application layer (Li et al., 2014; Kalia et al., 2019). For web-based applications, the response latencies of specific objects (e.g., those first painted or largest on the user’s screen) are highly correlated with several quality-of-experience metrics (E. Bocchi, L. De Cicco, and D. Rossi (2016); 22; 66; 5). As we elaborate in §2.1, to apply to a broad range of scenarios, we seek techniques to passively measure response latency while meeting the following additional requirements:
  1. (1)
    handle missing or encrypted transport headers;
  2. (2)
@@
  3. (3)
    generalize across transport dynamics;
  4. (4)
@@
Application-layer response latency is most closely related to the transport-layer round-trip time (RTT), whose passive measurement is widely studied, e.g.(Shawn Ostermann, ; Ghasemi et al., 2017; Sengupta et al., 2022; Veal et al., 2005; Jaiswal et al., 2004; Carra et al., 2010; Ding and Rabinovich, 2015; Jiang and Dovrolis, 2002). However, response latency is distinct from the transport-layer RTT, since the server may return a transport-layer acknowledgment well before the full application-layer response. As far as we are aware, no existing RTT measurement technique meets the additional requirements above, and we know of no passive techniques for directly measuring response latency (§2.2).
@@
Our first idea is to leverage the closed-loop nature of Internet protocols and applications. When application requests depend on the contents of prior responses—for example, a web object embeds other objects—the reception of a previous response generates the subsequent request. Flow-controlled applications (transports) mandate that new requests (packets) be transmitted only after prior responses (acknowledgments) arrive. Such closed-loop packet transmission behavior enables estimating the response latency by proxy: the vantage point can measure the time delay between a request and a subsequent request that was triggered by the reception of the response to the first request. We call the latter request a causally-triggered request, and the pair of requests a causal pair. Many latency-sensitive applications exhibit cross-request dependencies (e.g., web (Netravali et al., 2016), RPC (B. H. Sigelman, L. A. Barroso, M. Burrows, P. Stephenson, M. Plakal, D. Beaver, S. Jaspan, and C. Shanbhag (2010); S. Luo, H. Xu, C. Lu, K. Ye, G. Xu, L. Zhang, Y. Ding, J. He, and C. Xu (2021); Y. Gan, Y. Zhang, K. Hu, D. Cheng, Y. He, M. Pancholi, and C. Delimitrou (2019); 21)) and flow control (e.g., key-value store (Nishtala et al., 2013), web (98; 99)).
@@
It is not obvious how to identify causal pairs. At any given time, many concurrent requests may be in flight. Connection persistence and stream multiplexing is standardized and widely deployed in the HTTP protocols (48; 49; 47). Two consecutive requests arriving at a vantage point from the same connection do not necessarily form a causal pair.
@@
Our second key idea is to leverage the time gaps between packet arrivals to identify causal pairs. Once a client has sent a few requests, subsequent requests are typically blocked by dependencies or flow control. Hence, causally triggered requests arrive after a time gap longer than the inter-arrival times of the packets just prior. We call this the prominent packet gap assumption. By choosing an appropriate threshold for packet inter-arrival times, a vantage point can classify any packet arriving with a time gap exceeding this threshold as a causally-triggered request. The time interval between two consecutive causally-triggered requests provides one estimate of the response latency.
@@
But what should the time threshold be set to? How can we make this setting robust across deployments (e.g., wide-area vs. data center), network load, transport dynamics (e.g., TCP flavors), and applications (e.g.. web vs. key-value store)?
@@
Our third key idea is the following: While a single prominent packet gap is subject to (unknown) network, application, and server conditions, observing packet inter-arrivals over a period of time provides a more robust picture. In this paper, we leverage the probability distribution of inter-arrival times to identify prominent packet gaps by designing a lightweight construction to measure a coarse histogram of all packet inter-arrival times within a connection. We also devise a procedure to estimate the average response latency over the (configurable) time period when the distribution was measured.
@@
We call our algorithm, a synthesis of the three ideas above, Pirate. We show how to use measurements from Pirate in a feedback control loop to adapt a layer-4 load balancer. The load balancer may use direct server return (Patel et al., 2013) to avoid processing the responses. Since Pirate works under routing asymmetry, it is possible to make latency-aware decisions by unilaterally changing the load balancer, leaving clients, servers, application software, and the network unmodified.
Results. In §4, we evaluate Pirate under a realistic web workload derived from the Alexa top 100 web sites, with a web server whose CPU availability varies according to a real CPU utilization trace (Verma et al., 2015). Across all monitored responses, Pirate achieves a median relative error of 0.63% relative to the response latency measured at the client application, and gets to within $\pm$15% error 90% of the time. Our results also show that techniques for measuring transport-layer RTT do not faithfully measure application-layer response latency, even when they can leverage bidirectional traffic visibility. We integrated Pirate into Katran (86), an open-source layer-4 load balancer based on Maglev hashing (Eisenbud et al., 2016). Latency-aware Katran cuts the 99th percentile latency in our experimental setup by 37% on average across loads, relative to vanilla Katran.
@@
We outline caveats when interpreting our results in §5. A shorter version of this paper appeared previously at a workshop venue (citation elided for anonymity). In comparison, this paper incorporates new techniques to make estimation more robust, significantly expands evaluations of accuracy and overhead under realistic settings, and designs a feedback controller that operates on real-time measurements. This paper does not raise any ethical issues.
2. Motivation and Background
@@
Monitoring latency is a fundamental requirement for the design, maintenance, and optimization of interactive Internet services. This paper studies the problem of measuring response latency, which we define as the time between when an application client sends out the request and when it receives the last byte of the corresponding response at the application layer. The client may be a user device contacting a web server or an RPC client in one tier of a distributed multi-tiered application.
2.1. Our Goals
@@
We seek a passive measurement approach to continuously measure response latency at a vantage point that can lie outside the client and the server, along the path from the client to the server. To make it more broadly applicable, we also aim to achieve the following specific goals.
@@
G1. Handle missing or obscured transport headers. Encryption of application-layer payloads is ubiquitous on the web and in compute clusters (1; 19; 81; 70). Modern network-layer security goes one level further, obscuring transport-layer headers. Deployments of network-layer encrypted tunnels (e.g., WireGuard (126), IPSec (51)) have been trending upward (116). Even when transport is unencrypted, overlay networks and IP fragmentation can cause transport headers to be missing from packets. We desire a method that does not require visibility into the transport-layer headers.
@@
G2. Work under routing asymmetry. Nodes conducting passive measurement may be at locations that do not have access to traffic in both directions between the client and the server. The prevalence of routing asymmetry is well established in the wide-area Internet (Paxson, 1997) and also in data centers (Zhang et al., 2018). Further, many network deployments use asymmetric routing by design. For example, layer-4 load balancers employ direct server return (Patel et al., 2013; Eisenbud et al., 2016), bypassing the load balancer for response packets. BGP policies on the Internet (e.g., stub networks) may assign different ingress and egress border routers to a given connection (Katz-Bassett et al., 2010; Uhlig and Bonaventure, 2004).
@@
G3. Generalize across transport dynamics. The design and deployment of new transport algorithms for congestion control and loss recovery is an active and rapidly evolving area (Ferreira et al., 2024). We seek techniques that avoid relying on specific transport-protocol behaviors, and sidestep the need for new transport-protocols extensions.
@@
G4. Efficient enough to run in software middleboxes. Measurement and monitoring devices are frequently deployed as virtual network functions managed through a software-defined measurement infrastructure, e.g., (Palo Alto (2025); 117). We aim for techniques that impose low compute and memory overheads in software deployments, e.g., middleboxes running high-speed packet processing.
2.2. Prior Work and Its Applicability
In the academic literature, the problem most closely related to passively estimating response latency is the passive measurement of round-trip times (RTTs) at the transport layer.
@@ 2.2. Prior Work and Its Applicability
RTT estimation approaches. The basic idea of passive RTT estimation is to relate data packets to their corresponding transport-layer acknowledgments (ACKs) and measure the time difference between the pair. Many prior works on RTT estimation determine this relationship by associating the sequence numbers or timestamps in data packets with the corresponding information echoed in ACK packets (Shawn Ostermann, ; Ghasemi et al., 2017; Sengupta et al., 2022; Ding and Rabinovich, 2015; Veal et al., 2005; Jaiswal et al., 2004). Implementing this basic approach requires visibility into transport headers in the clear (G1) and seeing packets in both directions (G2).
@@
We are aware of only a small number of techniques that apply when the passive observer sees packet traffic in only one direction. An early approach by Jiang and Dovrolis (Jiang and Dovrolis, 2002) estimates RTT as the time interval between the SYN and the ACK packet of the 3-way handshake, and also between small packet bursts during the early rounds of slow start. These techniques are customized to the specific dynamics of the TCP protocol (G3) and do not measure connections beyond the initial phase. Another set of prior works estimate RTTs by computing frequency spectra from the time series of packet arrivals (Carra et al., 2010; Veal et al., 2005). Such computations require significant memory to hold packet timings (G4), and can be complex to tune (Ding and Rabinovich, 2015). Most recently, researchers have proposed the spin bit as a protocol extension (G3) to measure RTTs (De Vaere et al., 2018) passively. It has been incorporated as an optional mechanism in QUIC (115), but its deployment may be limited (Kunze et al., 2023).
@@
Response latency differs from RTT. A high RTT implies a high response latency, but a low RTT does not necessarily mean a low response latency. In principle, the response latency can differ significantly from the transport-layer RTT. The server can send a transport-layer ACK to the client before sending any response data. Further, responses, typically larger than requests, may span multiple packets, widening the gap between the transport-layer ACK and the last byte of the response. When the client and server use a protocol with pipelined requests (e.g., HTTP/1.1), multiple application requests may be transmitted in a single packet from the client to the server. The transport-layer ACK only corresponds to the first response.
Figure 1. Transport round-trip times (RTTs) do not match response latency, with clients running (a) TCP HTTP/1.1 and (b) UDP QUIC. See discussion in §$2.$
@@
Experimentally, we find that RTT and response latency differ even when there is only one small application-layer request in flight. We ran a web benchmark where a TCP client maintains a single outstanding request at a time with the server. Our full experimental setup is described in §4.1. On one client thread, Figure 1 (a) shows the time series of response latencies (marked req-to-res) against three techniques to measure transport RTT: (i) reading statistics from the kernel network stack through tcp-probe (113); (ii) estimating RTT from a packet trace (tcptrace (Shawn Ostermann, )); and (iii) syn-ack (Jiang and Dovrolis, 2002). Not only is the response latency different from the RTT, but it is also uncorrelated with it.
We ran a similar experiment to compare the RTT of a QUIC connection measured using the spin protocol extension (De Vaere et al., 2018) with the application-level response latency, measured with quic library lsquic (71) on a real web page. Figure 1 (b) shows that the two values diverge, especially when there are multiple RTT samples corresponding to a single response latency. The error bars show the minimum and maximum RTT reported through spin.
@@
Other related work. Given the emerging prevalence of encrypted transports, some recent work aims to passively infer quality of experience metrics specifically for web applications (Huet et al., 2020; Wehner et al., 2021). These metrics compose the response latencies of multiple objects, e.g., to estimate overall web experience (87), through machine learning and prediction. In contrast, we are concerned with continuously measuring the response latencies of individual objects throughout the lifetime of a connection. There is also significant prior work on passively measuring the latency of specific network segments, e.g.(Kompella et al., 2009; Lee et al., 2010; Ghabashneh et al., 2022; Canel et al., 2024) and on algorithms to estimate latencies over end-to-end network paths using previously measured delays, e.g.(Ng and Zhang, 2002; Madhyastha et al., 2006). Plenty of active approaches exist to measure end-to-end network latency characteristics, e.g.(Sommers et al., 2007; Yu et al., 2015; Sundaresan et al., 2017). This paper seeks continuous passive measurement techniques for end-to-end response latency.
@@
In summary, we believe that passive continuous measurement of response latencies under modern transport protocols and practical deployment constraints (§2.1) remains unsolved.
3. Pirate